# Perceptogram: Reconstructing Visual Percepts from EEG

### v2

@@ 3.1.1 From CLIP
NSD (Brain-Diffuser) Ozcelik and VanRullen, (2023) 0.254 0.356 0.942 0.962 0.872 0.915 0.775 0.423
NSD (MindEye) Scotti et al., (2023) $0.309$ $0.323$ $0.947$ $0.978$ $0.938$ $0.941$ $0.645$ $0.367$
@@
THINGS-MEG (BrainDecoding) Benchetrit et al., (2024) $0.088$ $0.333$ $0.747$ $0.855$ $0.712$ $0.804$ $0.576$
THINGS-MEG (EEGImageDecode) Li et al., 2024a - $0.340$ $0.613$ $0.672$ $0.619$ $0.603$ $0.651$
@@ 3.2.1 Lateral vs Medial negativity encodes various visual features
Fig. 9 shows one temporal slice displaying a spatial contrast between three semantic categories (animals, food, other), plotted on a 2-D topological scalp map. While these maps are inspired by the fMRI semantic maps from Huth et al., (2016), the temporal sensitivity of EEG permits the visualization of spatiotemporal maps that unfold in time, shown in Fig. 10. The semantic maps for individual subjects are shown in Fig. 28. Most subjects show the same spatial pattern seen in the grand-average from Fig. 10, with some individual differences in asymmetry and lateralization strength. The individual assymetries likely underlie the performance degradation we observe earlier from mirroring the electrodes.
@@ 4.1 Why does EEG visual reconstruction work?
In order to address the question of why we can do EEG visual reconstruction, specifically with a linear decoder, we investigate the representational similarity matrices (RSMs) in source (EEG), target (latent), and stimulus (image) space. We first consider the RSM of the EEG voltages, CLIP embeddings, and raw pixels for each of the 200 test images from the THINGS dataset (Fig. 13). The RSM plots of the EEG and CLIP embeddings show a clear organization by class — animals, food, and other. The pixels themselves appear to be structureless in this regime, suggesting that EEG and CLIP perform similar non-linear transformations from pixel to semantic space. As a result, their representational structures are aligned with respect to semantic class. Similar RSM comparison techniques have been widely used to evidence representational alignment between human brain and artificial neural network activations to the same stimuli (Allen et al., 2022b ; Tu et al., (2018); Yamins et al., (2013)).
This representational alignment between EEG and CLIP suggests that visual reconstruction via linear projection into CLIP latent space is made possible due to their shared structure. In our subsequent analysis, we compare this mapping to other image latent spaces: PCA, ICA, and VDVAE (see Appendix).
@@ 4.2 How does EEG visual reconstruction work?
We see that EEG shows similar representational relationships to those of the CLIP embeddings, which is why a simple linear decoder may be able to map the EEG to CLIP latents and reconstruct visual perceptions. The next question is, how does EEG encode such representational relationships? Our results describe specific features of the EEG that appear to change depending on the semantic content of the images. However, do they relate to plausible brain area activations? Answering these questions ensures that the relevant EEG features are not the results of behavioral artifacts or environmental noise.
@@
The semantic maps obtained in Fig. 10 open up a potentially new way to interpret the voltage 100-200 ms after stimulus onset. A negative polarity bias for faces at the occipito-temporal electrodes between 100-200ms is commonly known as the N170. Past studies have indicated its sensitivity to non-face categories, and particularly categories that are highly familiar (Rossion and Jacques,, 2008). Here, we provide another potential explanation to the N170 component amplitude – animacy. Animacy, being a common organizing principle in visual processing, involves activation in the ventral temporal cortex (Kriegeskorte et al.,, 2008).
@@ 4.2.2 Relating EEG to fMRI
In the fMRI dataset, categories such as “faces" and “humans from a distance" activate similar ventral lateral cortical area (see Fig. 34, 35) as “animals"; categories such as “room interiors", “urban scenes" (see Fig. 36, 37) and “food" do not. This presents the hypothesis that N170 is the EEG equivalent of the animacy axis.
@@ A.1 Dataset Details
Gifford et al., (2022) for EEG analysis, and the publicly available Natural Scenes Dataset (NSD) from Allen et al., 2022a for our fMRI analysis to validate findings from the EEG analysis.
@@ A.1.1 THINGS-EEG2
EEG data was collected while an image is presented for 100ms followed by a blank screen for 100ms before the next image. The image presentation order is pseudo-randomized across the entire image set. 10 subjects viewed the same 16740 images, of which the same 200 images are test images. We used the preprocessed version (https://osf.io/anp5v/), which has 17 posterior EEG channels compared to the 63 total channels in the raw dataset. The EEG was initially sampled at 1000Hz and down-sampled to 100Hz during the preprocessing. The only major filtering method applied during the preprocessing is Multi-Variate Noise Normalization (MVNN), which computes the covariance matrices of the EEG data (calculated for each time-point), and then averages them across image conditions and data partitions. The inverse of the resulting averaged covariance matrix is used to whiten the EEG data (independently for each session) (Gifford et al.,, 2022). Trials are extracted from $-0.2$s to 0.8s relative to the onset of the stimulus. Each training image is shown 4 times, and each test image is shown 80 times. We averaged all trials for the same image (within subject) to form the final dataset. At 100Hz sampling rate, $-0.2$s to 0.8 seconds corresponds to 100 samples. We discarded the first 20 samples which correspond to $-0.2$s to 0s, leaving 80 samples times 17 channels or 1360 dimensions per image per subject. The final dimensions of the training data for each subject are (16540 images, 1360 features) for the training set, and (200 images, 1360 features) for the test set.
@@ A.1.2 NSD
The fMRI data was collected while an image is presented for 3 seconds followed by a blank screen for 1 seconds before the next images. The images are taken from Microsoft’s COCO image database Lin et al., (2014) rather than the THINGS initiative image database. The image presentation order is pseudo-randomized across the entire image set: 982 images (used as the test set) are shared across 4 trial-complete subjects (1, 2, 5, 7) while each subject has 8859 images (used as the training set) exclusive to the particular subject and not shared across subjects. Each image is presented 3 times to enhance the signal-to-noise ratio, and the presentation order of which is controlled.
@@ A.3 Evaluation Metrics
We used the same performance metrics (see Table 1) as in Ozcelik and VanRullen, (2023), which has been used in other followup studies such as MindEye Scotti et al., (2023). The 8 metrics we used are Pixel Correlation (PixCorr), Structural Similarity (SSIM), AlexNet layer 2 and 5 outputs pairwise correlations, InceptionNet output pairwise correlation, CLIP ViT output pairwise correlation, EfficientNet output distance, and SwAV output distance. PixCorr and SSIM involve comparing the reconstructed image with the ground-truth (GT) test image. PixCorr is a low-level (pixel) measure that involves vectorizing the reconstructed and GT images and computing the correlation coefficient between the resulting vectors. SSIM is a measure developed by Wang et al. 2004 that computes a match value between GT and reconstructed images as a function of overall luminance match, overall contrast match, and a “structural” match which is defined by normalizing each image by its mean and standard deviation. We should note that this measure was designed for comparing images with minor distortions and does not seem as reliable for images that are not close matches as currently obtained with image reconstruction methods. One way to see this is to observe that the SSIM measure does not seem much affected by the duration of EEG window over 50ms, unlike the other measures that show performance improvements from 100ms through 400ms.

### v3

@@ 3.1.1 From CLIP
NSD (Brain-Diffuser) Ozcelik_VanRullen_2023 0.254 0.356 0.942 0.962 0.872 0.915 0.775 0.423
NSD (MindEye) Scotti_Banerjee_Goode_Shabalin_Nguyen_Cohen_Dempster_Verlinde_Yundler_Weisberg_et_al_2023 $0.309$ $0.323$ $0.947$ $0.978$ $0.938$ $0.941$ $0.645$ $0.367$
@@
THINGS-MEG (BrainDecoding) Benchetrit_Banville_King $0.088$ $0.333$ $0.747$ $0.855$ $0.712$ $0.804$ - $0.576$
THINGS-MEG (EEGImageDecode) Li_Wei_Li_Zou_Liu_2024 - $0.340$ $0.613$ $0.672$ $0.619$ $0.603$ - $0.651$
@@ 3.1.4 Time-Swapping
3.2 Decoding performance
We trained a linear regression with no bias term from zero-mean EEG data to CLIP embeddings. We also trained a linear network with single layer from EEG dimension to CLIP dimension directly with no bias using MSE loss. The predicted CLIP embeddings were divided by their own L2-norm and multiplied by the mean training CLIP L2-norm. Both model were evaluated using the retrieval metric based on recall@1 and recall@5.
Table 3: Retrieval performance of EEG-to-CLIP decoding model.
Method Recall@1 (%) Recall@5 (%)
Linear Regression 25.0 50.0
Linear network 24.5 49.5
3.3 Encoding Weight Reconstructions Visualizes Presumptive Neural Preference
Figure 8: Encoding Weight Reconstructions Visualizes Presumptive Neural Preference (Subject 1). Each row represents an electrode in the visual cortical area and each column represents a 10ms time bin. Each individual image is a visual reconstruction of the linear encoding weight from CLIP to that particular electrode-time-bin, representing the CLIP visual features that would maximize the voltage at that point. But because here the linear weights are multiplied by -1, in this example it would mean the visual features that would minimize the voltage at that point. This would allow for the interpretation of negative ERP components. For example, there are 2 clusters of face images for electrode P3, P7, P5, and PO7 at around 160-200ms, so the face CLIP features would minimize the voltage for those electrodes around that time. The lateral location of these electrode and timing matches the N170 ERP component, which is a negative going ERP component around 170ms sensitive to faces. Because assigning meaning to either polarity involves somewhat of a subjective decision, we show in the next figure the positive weight reconstructions just as a comparison.
Figure 9: The positive weight version of Presumptive Neural Preference (Subject 1). Where the face cluster used to be for the negative weight reconstructions are now replaced by some kind of scenes here.
To directly visualize what visual features drive EEG responses, we reconstructed preferred stimuli from the CLIP-to-EEG encoding weights using unCLIP. This approach treats each weight vector as a CLIP embedding, revealing the "ideal stimulus" that would maximally activate each electrode according to our learned linear model.
The spatiotemporal reconstruction maps for subject 1 showed a high spatial frequency cluster for the medial channels such as O1, Oz, and O2 around 100-180ms, and a face cluster for the lateral channels such as P3, P7, P5, PO7 (in this case, all on the left-side) around 160-200ms. In this case, the weights are inverted (multiplied by -1), it is to show the CLIP features that give rise to negative voltages. The positive version of the weights are also included as a comparison.
@@ 3.4.1 Lateral vs Medial negativity encodes various visual features
Fig. 11 shows one temporal slice displaying a spatial contrast between three semantic categories (animals, food, other), plotted on a 2-D topological scalp map. While these maps are inspired by the fMRI semantic maps from Huth2016, the temporal sensitivity of EEG permits the visualization of spatiotemporal maps that unfold in time, shown in Fig. 12. The semantic maps for individual subjects are shown in Fig. 31. Most subjects show the same spatial pattern seen in the grand-average from Fig. 12, with some individual differences in asymmetry and lateralization strength. The individual assymetries likely underlie the performance degradation we observe earlier from mirroring the electrodes.
@@ 4 Discussion
We present a unified framework that not only enables image reconstruction from EEG using a simple linear decoder, but also isolates the latent-filtered EEG patterns that support visual perception. Our findings challenge the prevailing assumption that high-fidelity visual decoding from EEG requires deep, non-linear models. Instead, we demonstrate that EEG representations are inherently compatible with CLIP’s latent structure, enabling direct linear mapping for high-quality image reconstructions. Alignment with spatial patterns derived from fMRI further validates the spatial organization captured by our EEG-derived features, while also highlighting the unique temporal dynamics accessible through EEG. In the following two sections, we discuss (a) Why a simple linear mapping is sufficient to achieve the reported reconstruction performance, and (b) what our perturbation experiments and latent-filtered EEG patterns reveal about the spatial and temporal organization of visual representations in the EEG.
@@ 4.1 Why does EEG visual reconstruction work?
Figure 16: Subject-wise CLIP–EEG RSM similarity vs. reconstruction performance (Spearman $r=0.579$).
Unlike deep learning methods that rely on multi-layered feature extraction, Perceptogram linearly maps EEG responses to CLIP latents.
To understand why EEG visual reconstruction is possible using a simple linear decoder, we examine the representational similarity matrices (RSMs) across EEG, CLIP, and pixel space (Fig. 15). While pixel-based representations show little structure, both EEG and CLIP embeddings organize stimuli into clear semantic clusters (e.g., animals vs. food). This shared structure (average Spearman correlation between EEG and CLIP test RSMs is .22) suggests that both the brain and CLIP perform nonlinear transformations from raw visual input to a semantically structured latent space, enabling their linear alignment (see Fig. 16). Alignment has also been shown in previous comparisons between neural and deep network activations Allen_St-Yves_Wu_Breedlove_Prince_Dowdle_Nau_Caron_Pestilli_Charest_etal._2022; tu2018relating; yamins2013hierarchical.
@@ 4.2 How does EEG visual reconstruction work?
Our perturbation experiments and latent-filtered EEG patterns reveal key insights into how visual representations are encoded in EEG signals, both spatially and temporally. In line with prior work, our time-swapping experiments suggest that visual information is distributed across distinct temporal windows. Specifically, low-level visual features such as fur color are encoded earlier, while higher-level semantic content—such as animacy—is encoded later in the EEG time course. Spatially, swapping only the O1 and O2 electrodes can flip reconstructions across the vertical midline, consistent with the known contralateral organization of early visual processing. In contrast, mirroring all electrodes induces broader changes, affecting both visual and semantic attributes in the reconstruction. Notably, some reconstructions originally depicting animate objects appear inanimate after mirroring, and vice versa.
This finding, together with the subject-specific CLIP-filtered EEG spatial patterns associated with animals, provides converging evidence that animacy may be encoded asymmetrically and laterally at the individual level. In contrast, EEG responses to food appear more medially organized and less sensitive to mirroring. This medial-lateral distinction also corresponds to other visual features (e.g., texture, brightness), with individual differences in lateralization, as illustrated in Fig. 31.
@@
The semantic maps obtained in Fig. 12 open up a potentially new way to interpret the voltage 100-200 ms after stimulus onset. A negative polarity bias for faces at the occipito-temporal electrodes between 100-200ms is commonly known as the N170. Past studies have indicated its sensitivity to non-face categories, and particularly categories that are highly familiar (Rossion2008). Here, we provide another potential explanation to the N170 component amplitude – animacy. Animacy, being a common organizing principle in visual processing, involves activation in the ventral temporal cortex (kriegeskorteRepresentationalSimilarityAnalysis2008).
@@ 4.2.2 Relating EEG to fMRI
In the fMRI dataset, categories such as “faces" and “humans from a distance" activate similar ventral lateral cortical area (see Fig. 37, 38) as “animals"; categories such as “room interiors", “urban scenes" (see Fig. 39, 40) and “food" do not. This presents the hypothesis that N170 is the EEG equivalent of the animacy axis.
@@ A.1 Dataset Details
Gifford_Dwivedi_Roig_Cichy_2022 for EEG analysis, and the publicly available Natural Scenes Dataset (NSD) from allen2022massive for our fMRI analysis to validate findings from the EEG analysis.
@@ A.1.1 THINGS-EEG2
EEG data was collected while an image is presented for 100ms followed by a blank screen for 100ms before the next image. The image presentation order is pseudo-randomized across the entire image set. 10 subjects viewed the same 16740 images, of which the same 200 images are test images. We used the preprocessed version (https://osf.io/anp5v/), which has 17 posterior EEG channels compared to the 63 total channels in the raw dataset. The EEG was initially sampled at 1000Hz and down-sampled to 100Hz during the preprocessing. The only major filtering method applied during the preprocessing is Multi-Variate Noise Normalization (MVNN), which computes the covariance matrices of the EEG data (calculated for each time-point), and then averages them across image conditions and data partitions. The inverse of the resulting averaged covariance matrix is used to whiten the EEG data (independently for each session) (Gifford_Dwivedi_Roig_Cichy_2022). Trials are extracted from 0.2-0.2s to 0.8s relative to the onset of the stimulus. Each training image is shown 4 times, and each test image is shown 80 times. We averaged all trials for the same image (within subject) to form the final dataset. At 100Hz sampling rate, 0.2-0.2s to 0.8 seconds corresponds to 100 samples. We discarded the first 20 samples which correspond to 0.2-0.2s to 0s, leaving 80 samples times 17 channels or 1360 dimensions per image per subject. The final dimensions of the training data for each subject are (16540 images, 1360 features) for the training set, and (200 images, 1360 features) for the test set.
@@ A.1.2 NSD
The fMRI data was collected while an image is presented for 3 seconds followed by a blank screen for 1 seconds before the next images. The images are taken from Microsoft’s COCO image database coco_dataset_2014 rather than the THINGS initiative image database. The image presentation order is pseudo-randomized across the entire image set: 982 images (used as the test set) are shared across 4 trial-complete subjects (1, 2, 5, 7) while each subject has 8859 images (used as the training set) exclusive to the particular subject and not shared across subjects. Each image is presented 3 times to enhance the signal-to-noise ratio, and the presentation order of which is controlled.
@@ A.3 Evaluation Metrics
We used the same performance metrics (see Table 2) as in Ozcelik_VanRullen_2023, which has been used in other followup studies such as MindEye Scotti_Banerjee_Goode_Shabalin_Nguyen_Cohen_Dempster_Verlinde_Yundler_Weisberg_et_al_2023. The 8 metrics we used are Pixel Correlation (PixCorr), Structural Similarity (SSIM), AlexNet layer 2 and 5 outputs pairwise correlations, InceptionNet output pairwise correlation, CLIP ViT output pairwise correlation, EfficientNet output distance, and SwAV output distance. PixCorr and SSIM involve comparing the reconstructed image with the ground-truth (GT) test image. PixCorr is a low-level (pixel) measure that involves vectorizing the reconstructed and GT images and computing the correlation coefficient between the resulting vectors. SSIM is a measure developed by Wang et al. 2004 that computes a match value between GT and reconstructed images as a function of overall luminance match, overall contrast match, and a “structural” match which is defined by normalizing each image by its mean and standard deviation. We should note that this measure was designed for comparing images with minor distortions and does not seem as reliable for images that are not close matches as currently obtained with image reconstruction methods. One way to see this is to observe that the SSIM measure does not seem much affected by the duration of EEG window over 50ms, unlike the other measures that show performance improvements from 100ms through 400ms.